# Gaussian is All You Need: A Unified Framework for Solving Inverse Problems via Diffusion Posterior Sampling

### v2

@@ 4.1 Experiment on mixture of gaussians
We performed image restoration experiments using the FFHQ [19] and ImageNet [13] datasets. Our experiments include image super-resolution, random inpainting, and Gaussian and motion deblurring tasks. We compare our method with related diffusion-based approaches such as $\Pi$GDM [37], DDRM [21], DiffPIR [48], DPS [9], DMPS [24] , MCG [8], DDS[10], DDNM[41] and ReSample[35]. Additionally, we compare it with ADMM for Total Variation (TV) regularization [32, 15] and PnP-ADMM [6] using a DnCNN [46] denoiser. For each of these methods, we used the publicly available source code with the reported hyper-parameters. Further details of the experimental setup, hyper-parameters for our method, as well as for all the comparison methods, are described in Appendix B.
@@
Inverse problems. The first inverse problem we consider is image super-resolution. The task here is to recover an image blurred using a $9\times 9$ Gaussian blur kernel with standard deviation $3.0$ followed by a ×$\times 4$ down-sampling using a decimation matrix. The down-sampling operator is a decimation matrix that samples the first pixel in every non-overlapping $4\times 4$ block of our image. For Gaussian deblurring, we use a $61\times 61$ kernel with standard deviation $3.0$ and for motion deblurring we randomly generate a kernel [^1] https://github.com/LeviBorodenko/motionblur with size $61\times 61$ with intensity of $0.5$. Finally, we consider random image inpainting, where we remove pixels randomly in all color channels. Each pixel can be removed with a uniform probability in the range $[0.7,0.8]$. For noisy experiments, we apply an additive Gaussian noise with standard deviation $\bm{\sigma}_{n}=0.05$.
SR (×$\times 4$) Deblur (Gaussian) Deblur (motion)
@@
Table 2: Quantitative metrics for noiseless inverse problems on the ImageNet dataset. Our proposed method is approximately twice as computationally efficient as $\Pi$GDM and achieves the best performance in deblurring tasks.
Results. We present quantitative metrics on the FFHQ dataset in Tables 1 and 3. In both motion and Gaussian deblurring experiments, our methods with 1000 and 100 number of function evaluations (NFEs) outperformed all competing methods across several metrics. With the exception of the FID scores, our methods consistently delivered either the best or second-best performance. The only two exceptions are: DPS achieved the top FID score for Gaussian deblurring, while DiffPIR obtained the best FID score for motion deblurring. These results demonstrate the effectiveness of our approaches in both deblurring tasks. For 4×4$\times$4 × super-resolution experiments, our method demonstrates strong performance, achieving a PSNR of 26.34 dB and an SSIM of 0.751. While DDRM achieves a marginally higher PSNR (+0.13 dB), CoDPS demonstrates superior perceptual quality. When compared to DMPS, although CoDPS shows lower perceptual quality, its performance remains highly competitive with comparable PSNR, SSIM and LPIPS metrics.
@@
DPS [9] 29.12 0.852 27.81 0.195
ReSample [35] 27.78 0.795 54.52 0.274
DDS [10] 27.01 0.743 46.99 0.325
DDNM [41] 28.69 0.840 36.84 0.218
DMPS [24] 21.33 0.573 112.30 0.573
DDRM [21] 25.26 0.761 61.57 0.295
MCG [8] 26.59 0.778 31.94 0.231
PnP-ADMM [6] 23.77 0.727 57.00 0.360
ADMM-TV [15] 20.50 0.623 141.45 0.568
@@
We designed an additional noiseless experiment to compare our method with $\Pi$GDM and DDNM, and we report the results in Table 2. From the noisy experiments, we observed that both methods $\Pi$GDM and DDNM+ did not perform well. As a result, we designed an experiment that closely matches the setup in the original papers. We used the ImageNet dataset on three tasks: super-resolution, motion deblurring, and Gaussian deblurring. Our method outperformed $\Pi$GDM and DDNM by a large margin in the Gaussian deblurring experiment, while $\Pi$GDM had the upper hand in the super-resolution experiment. For motion deblurring, our method achieved the best metrics in terms of SSIM, FID, and LPIPS, while $\Pi$GDM achieved the best PSNR scores. We did not include motion deblurring results because the current implementation of DDNM does not support it. Overall, our method produced consistent and improved results in both noiseless and noisy problems.
Table 4: Comparison of efficiency metrics for various diffusion-based solvers on SR (×4)($\times$ 4)( × 4 ) using the FFHQ dataset, highlighting the number of function evaluations (NFE), wall-clock time in seconds and peak GPU memory usage when performing inference on a single $256\times 256$ RGB image.
Method NFE Wall-clock time (Secs/img) $\downarrow$ GPU Memory Req (in GB) $\downarrow$
ReSample 100 230.4 4.32
DDNM 1000 45.73 4.73
DPS 1000 146.8 9.35
DMPS 1000 63.9 4.33
$\Pi$GDM 1000 81.4 9.90
@@
Algorithm efficiency. Our method is both memory and time efficient compared to related methods such as DPS and $\Pi$GDM. Both these methods need to compute gradients through the diffusion networks, which can increase the cost per NFE by a factor of $2$-$3$ as discussed in [37]. In Table 4, we report the average wall-clock time per image in seconds and the peak GPU memory requirement during inference for each method. For a consistent NFE of 1000, our method, CoDPS, is approximately 2×2$\times$2 × faster on average compared to DPS, achieving a wall-clock time of 42.7 seconds per image. Additionally, CoDPS shows competitive memory efficiency to DMPS and ReSample, requiring only 4.73 GB of GPU memory. However, the runtime of ReSample is substantially worse compared to our method. ReSample uses a latent diffusion-based model, where the latent image size is 64. This provides some advantage in terms of computation, as the diffusion process occurs over a smaller size ($64\times 64$) compared to the pixel space ($256\times 256$). However, the decoding and ReSampling stages can be computationally expensive, which resulted in a significantly longer runtime compared to our method. Similar to DMPS, our method avoids gradient computations through the diffusion networks, leading to improved memory efficiency. Moreover, CoDPS significantly outperforms both DPS and $\Pi$GDM in memory efficiency. These findings show the effectiveness of CoDPS in optimizing computational resources and make it suitable for resource constrained settings.
@@ A.2 Proof of theorem 1
We observe that this exponent is quadratic in $x$ and $y$ and we express it as a bi-linear form using a matrix $\bm{M}$. Since the exponent in (18) does not any terms that contain $x^{2}y^{2}$,$xy^{2}$,and $x^{2}y$, we must have $m_{21}=m_{12}=m_{22}=0$. We can also easily see that $m_{20}=\dfrac{-1}{2(1-\rho^{2})\sigma^{2}_{x}}$ , $m_{02}=\dfrac{-1}{2(1-\rho^{2})\sigma^{2}_{y}}$, $m_{11}=\dfrac{-\rho}{(1-\rho^{2})\sigma_{x}\sigma_{y}}$, and $m_{11}^{2}=4\rho^{2}m_{20}m_{10}$. Since $10<\rho^{2}<$ 1, we have $m_{11}^{2}<4m_{20}m_{10}$. ∎
@@
where $r$ is an integer between 00 and $n$, $\varphi_{i}(x)$ and $\Psi_{j}(y)$ are arbitrary independent functions, while the each $a_{ij}$ and $b_{ij}$ are constants that satisfy
@@
[…]
@@ A.3 Derivation for conditional score
[…]
@@ A.4 Efficient covariance inversion
When $k=l$, the exponent becomes zero, resulting in the terms summing up to $m$. If $l-k<m$ and $l\neq m$, we can apply the geometric sum formula as
@@
If $l-k>m$, due to the periodicity of the $N$th roots of unity, the values will be repeating. Thus we obtain $d$ blocks $m\times m$ identity matrices stacked horizontally.
@@ B.1 Experimental setup
The first inverse problem we consider is image super resolution. The task here is to recover an image blurred using a $9\times 9$ Gaussian blur kernel with standard deviation $3.0$ followed by a ×$\times 4$ down-sampling using a decimation matrix. The down-sampling operator is a decimation matrix that samples the first pixel in every non-overlapping $4\times 4$ block of our image. For Gaussian deblurring, we use a $61\times 61$ kernel with standard deviation $3.0$ and for motion delubrring we randomly generate a kernel [^2] https://github.com/LeviBorodenko/motionblur with size $61\times 61$ with intensity of $0.5$. Finally, we consider random image inpainting, where we remove pixels randomly in all color channels. Each pixel can be removed with a uniform probability in the range $[0.7,0.8]$. For noisy experiments, we apply an additive Gaussian noise with standard deviation $\bm{\sigma}_{n}=0.05$.
@@ B.3 Implementation details
CoDPS. Our algorithm has two main parameters the prior data covariance $\sigma_{x(0)}^{2}$ and the gradient scalar term $\{\zeta_{t}\}$ as shown in Algorithms 1 and 2. For the DDIM sampling, we set $\eta=1$ and NFE$=100$ for all experiments. We have to main hyper-parameters that we tune for various problems. These are $\{\zeta_{t}\}_{t=1}^{t=T}$ and $\sigma^{2}_{\bm{x(0)}}$. For $\{\zeta_{t}\}_{t=1}^{T}$, we start with a fixed value and then reduce it once based on a condition related to $t$. We report the values used in the main paper for each inverse problems in Tables 5 and 6. We also report a variant of our method that uses DDPM sampling in Algorithm 2. Both DDIM and DDPM algorithms requires the total number of diffusion steps $N$, a measurement vector $\bm{y}$, scaling parameters $\{\zeta_{t}\}_{t=1}^{N}$, noise standard deviation ${\{\tilde{\sigma}_{t}\}_{t=1}^{N}}$, and prior variance hyper-parameter $\sigma_{0}^{2}$ and outputs the final reconstructed image $\hat{\bm{x}}_{0}$.
@@
×$\times 4$ SR Gaussian Deblur Motion Deblur Inpainting
@@
[…]
ImageNet DDIM $\zeta_{t}$$\begin{cases}5\text{e-2}&\text{if }t>20\\ 1\text{e-2}&\text{otherwise }\end{cases}$$\begin{cases}4\text{e-2}&\text{if }t>15\\ 8\text{e-3}&\text{otherwise }\end{cases}$$\begin{cases}4\text{e-2}&\text{if }t>25\\ 1\text{e-2}&\text{otherwise }\end{cases}$
[…]
×$\times 4$ SR Gaussian Deblur Motion Deblur
[…]
Algorithm 2 CoDPS(DDPM)
0:  $N$, $\bm{y}$, $\zeta_{t}\}_{t=1}^{N},{\{\tilde{\sigma}_{t}\}_{t=1}^{N}}$, $\sigma_{x_{0}}^{2}$
1:  $\bm{x}_{N}\sim\mathcal{N}(\bm{0},\bm{I})$
2:  for $t=N-1$ to 00 do
3:     $\hat{\bm{s}}\leftarrow\bm{s}_{\theta}(\bm{x}_{t},t)$
4:     $\hat{\bm{x}}_{0}\leftarrow\dfrac{1}{\sqrt{\bar{\alpha}_{t}}}(\bm{x}_{t}+(1-\bar{\alpha}_{t})\hat{\bm{s}})$
@@
$\Pi$GDM. We obtained the implementation of $\Pi$GDM from [37] for the noiseless inverse problems. We used the pseudo-code from the paper to implement the noisy version of the algorithm. We used DDIM sampling and set NFE to 100 for all experiments.
@@
DDS. We used the publicly available version of DDS [10]. We included our forward operator implementations, since the original code only supported medical imaging tasks. We tuned the regularizer parameter $\gamma$ between range $0.5-2$ using grid search to obtain the best result for each problem.
@@
ReSample. We used the publicly available implementation of ReSample [35]. For each task, we used the corresponding configuration files. We modified the noise-level configuration to $\sigma_{y}=0.05$. For each problem, we tuned the scale parameters within the range 0.10.10.10.11111.
@@ Appendix C Additional results and figures
Learning the data prior. In this experiment, we limit the number of mixtures to $k=3$. We build our dataset by generating samples from an $n$-dimensional GMM model, where $n$ is a randomly selected integer between $3$ and $15$. We randomly set the means of mixtures using random $n$ dimension vectors selected uniformly form the range $[-3,3]$ and the covariance matrices using an $n\times n$ symmetric positive definite matrix whose entries are sampled uniformly from the range $[0,1]$. Now that we have a data at hand the first task it to learn its distribution using a diffusion model. We use an MLP-based model similar to the one used for toy problems in [34]. Our diffusion model consists of 4 MLP blocks, each with an input and output feature size of 64. Additionally, the model has input and output layers to map features between the data dimension $n$ and 64.
@@ C.2.1 Hyperparameter ablation study
To assess the sensitivity of our proposed methods to hyperparameters, we conducted an ablation study based on the reviewer’s suggestion. Specifically, we used the $\sigma_{0}^{2}$ and gradient scaling schedule ${\zeta_{t}}$ tuned for the super-resolution task (SR ×$\times 4$), and then applied the same hyperparameters directly to the Gaussian deblurring, motion deblurring, and inpainting tasks on the FFHQ dataset.
@@
Table 7: Ablation study using hyperparameters tuned for SR (×$\times 4$) and applied directly to other tasks. Results show that CoDPS remains competitive without per-task tuning on two new tasks.
@@
SR (×$\times 4$) Deblur (Gaussian) Deblur (motion)
@@
$\Pi$GDM [37] 100 21.23 0.482 88.93 0.495 18.29 0.341 135.86 0.610 23.89 0.663 52.15 0.373
DPS [9] 1000 21.70 0.562 45.20 0.381 21.97 0.706 62.72 0.444 20.55 0.634 56.08 0.389
DiffPIR [48] 100 20.58 0.430 110.69 0.560 21.62 0.504 58.61 0.433 24.61 0.661 46.72 0.359
DDRM [21] 20 23.64 0.624 67.69 0.423 22.73 0.705 63.02 0.427 - - - -
MCG [8] 1000 13.39 0.227 144.5 0.637 16.32 0.441 95.04 0.550 5.89 0.037 186.9 0.758
PnP-ADMM [6] - 20.43 0.538 159.58 0.592 20.77 0.545 105.96 0.502 21.02 0.563 108.53 0.500
ADMM-TV [15] - 18.40 0.490 191.97 0.600 21.52 0.594 108.10 0.495 20.39 0.557 148.80 0.524
@@
Table 9: Performance metrics of inverse problems on ImageNet 256$\times$256-1k validation dataset.
@@
DPS [9] 25.02 0.750 32.21 0.269
DiffPIR [48] - - - -
DDRM [21] 22.06 0.593 84.44 0.437
MCG [8] 22.62 0.608 30.51 0.316
PnP-ADMM [6] 21.41 0.641 65.55 0.403
ADMM-TV 18.94 0.546 138.40 0.540

### v3

@@ 4.1 Experiment on mixture of gaussians
[…]
@@
Inverse problems. The first inverse problem we consider is image super-resolution. The task here is to recover an image blurred using a $9\times 9$ Gaussian blur kernel with standard deviation $3.0$ followed by a $\times 4$ down-sampling using a decimation matrix. The down-sampling operator is a decimation matrix that samples the first pixel in every non-overlapping $4\times 4$ block of our image. For Gaussian deblurring, we use a $61\times 61$ kernel with standard deviation $3.0$ and for motion deblurring we randomly generate a kernel [^1] https://github.com/LeviBorodenko/motionblur with size $61\times 61$ with intensity of $0.5$. Finally, we consider random image inpainting, where we remove pixels randomly in all color channels. Each pixel can be removed with a uniform probability in the range $[0.7,0.8]$. For noisy experiments, we apply an additive Gaussian noise with standard deviation $\bm{\sigma}_{n}=0.05$.
SR ($\times 4$) Deblur (Gaussian) Deblur (motion)
@@
[…]
Results. We present quantitative metrics on the FFHQ dataset in Tables 1 and 3. In both motion and Gaussian deblurring experiments, our methods with 1000 and 100 number of function evaluations (NFEs) outperformed all competing methods across several metrics. With the exception of the FID scores, our methods consistently delivered either the best or second-best performance. The only two exceptions are: DPS achieved the top FID score for Gaussian deblurring, while DiffPIR obtained the best FID score for motion deblurring. These results demonstrate the effectiveness of our approaches in both deblurring tasks. For $4\times$ super-resolution experiments, our method demonstrates strong performance, achieving a PSNR of 26.34 dB and an SSIM of 0.751. While DDRM achieves a marginally higher PSNR (+0.13 dB), CoDPS demonstrates superior perceptual quality. When compared to DMPS, although CoDPS shows lower perceptual quality, its performance remains highly competitive with comparable PSNR, SSIM and LPIPS metrics.
@@
DPS [7]
29.12 0.852 27.81 0.195
ReSample [16]
27.78 0.795 54.52 0.274
DDS [9]
27.01 0.743 46.99 0.325
DDNM [15]
28.69 0.840 36.84 0.218
DMPS [14]
21.33 0.573 112.30 0.573
DDRM [5]
25.26 0.761 61.57 0.295
MCG [20]
26.59 0.778 31.94 0.231
PnP-ADMM [33]
23.77 0.727 57.00 0.360
ADMM-TV [34]
20.50 0.623 141.45 0.568
@@
[…]
Table 4: Comparison of efficiency metrics for various diffusion-based solvers on SR $(\times 4)$ using the FFHQ dataset, highlighting the number of function evaluations (NFE), wall-clock time in seconds and peak GPU memory usage when performing inference on a single $256\times 256$ RGB image.
Method NFE
Wall-clock time
(Secs/img)
$\downarrow$
GPU Memory Req
(in GB) $\downarrow$
RED-DIFF
1000 101.8 9.44
DDS
1000 45.6 4.73
DiffPIR
1000 45.7 4.22
MCG
1000 84.0 9.56
DDRM
1000 54.7 4.73
ReSample
100 230.4 4.32
DDNM
1000 45.73 4.73
DPS
1000 146.8 9.35
DMPS
1000 63.9 4.33
[…]
1000 81.4 9.90
@@
[…]
@@ A.2 Proof of theorem 1
We observe that this exponent is quadratic in $x$ and $y$ and we express it as a bi-linear form using a matrix $\bm{M}$. Since the exponent in (18) does not any terms that contain $x^{2}y^{2}$$xy^{2}$,and $x^{2}y$, we must have $m_{21}=m_{12}=m_{22}=0$. We can also easily see that $m_{20}=\dfrac{-1}{2(1-\rho^{2})\sigma^{2}_{x}}$ , $m_{02}=\dfrac{-1}{2(1-\rho^{2})\sigma^{2}_{y}}$, $m_{11}=\dfrac{-\rho}{(1-\rho^{2})\sigma_{x}\sigma_{y}}$, and $m_{11}^{2}=4\rho^{2}m_{20}m_{10}$. Since $0<\rho^{2}<1$, we have $m_{11}^{2}<4m_{20}m_{10}$. ∎
@@
where $r$ is an integer between 0 and $n$, $\varphi_{i}(x)$ and $\Psi_{j}(y)$ are arbitrary independent functions, while the each $a_{ij}$ and $b_{ij}$ are constants that satisfy
@@
$\begin{aligned} &\sigma_{X|Y}^{2}[2\sigma_{Y|X}^{2}u(x)-\mu_{Y|X}^{2}]+\sigma_{Y|X}^{2}[\mu_{X|Y}^{2}-2\sigma_{X|Y}^{2}v(y)]\\ &-y^{2}\mu_{X|Y}^{2}+x^{2}\sigma_{Y|X}^{2}+2(\sigma_{X|Y}^{2}\mu_{Y|X}y-\sigma_{Y|X}^{2}\mu_{X|Y}x)=0\end{aligned}.$ (27)
@@ A.3 Derivation for conditional score
$p(\bm{y}|\bm{x}_{t})$ $=\int p(\bm{y}|\bm{z},\bm{x}_{t})p(\bm{z}|\bm{x}_{t})d\bm{z}=\int p(\bm{y}|\bm{z})p(\bm{z}|\bm{x}_{t})d\bm{z}$
$=\int d\bm{z}\underbrace{\frac{1}{C_{1}}\exp\left(-\frac{\|\bm{y}-\bm{z}\|^{2}_{2})}{2\sigma_{n}^{2}}\right)}_{\mathcal{F}(\bm{x})}$
@@ A.4 Efficient covariance inversion
When $k=l$, the exponent becomes zero, resulting in the terms summing up to $m$. If lk<ml-k<m and $l\neq m$, we can apply the geometric sum formula as
@@
If lk>ml-k>m, due to the periodicity of the $N$th roots of unity, the values will be repeating. Thus we obtain $d$ blocks $m\times m$ identity matrices stacked horizontally.
@@ B.1 Experimental setup
The first inverse problem we consider is image super resolution. The task here is to recover an image blurred using a $9\times 9$ Gaussian blur kernel with standard deviation $3.0$ followed by a $\times 4$ down-sampling using a decimation matrix. The down-sampling operator is a decimation matrix that samples the first pixel in every non-overlapping $4\times 4$ block of our image. For Gaussian deblurring, we use a $61\times 61$ kernel with standard deviation $3.0$ and for motion delubrring we randomly generate a kernel [^2] https://github.com/LeviBorodenko/motionblur with size $61\times 61$ with intensity of $0.5$. Finally, we consider random image inpainting, where we remove pixels randomly in all color channels. Each pixel can be removed with a uniform probability in the range $[0.7,0.8]$. For noisy experiments, we apply an additive Gaussian noise with standard deviation $\bm{\sigma}_{n}=0.05$.
@@ B.3 Implementation details
CoDPS. Our algorithm has two main parameters the prior data covariance $\sigma_{x(0)}^{2}$ and the gradient scalar term $\{\zeta_{t}\}$ as shown in Algorithms 1 and 2. For the DDIM sampling, we set $\eta=1$ and NFE=100=100 for all experiments. We have to main hyper-parameters that we tune for various problems. These are $\{\zeta_{t}\}_{t=1}^{t=T}$ and $\sigma^{2}_{\bm{x(0)}}$. For $\{\zeta_{t}\}_{t=1}^{T}$, we start with a fixed value and then reduce it once based on a condition related to $t$. We report the values used in the main paper for each inverse problems in Tables 5 and 6. We also report a variant of our method that uses DDPM sampling in Algorithm 2. Both DDIM and DDPM algorithms requires the total number of diffusion steps $N$, a measurement vector $\bm{y}$, scaling parameters $\{\zeta_{t}\}_{t=1}^{N}$, noise standard deviation $\{\tilde{\sigma}_{t}\}_{t=1}^{N}$}, and prior variance hyper-parameter $\sigma_{0}^{2}$ and outputs the final reconstructed image $\hat{\bm{x}}_{0}$.
@@
$\times 4$ SR Gaussian Deblur Motion Deblur Inpainting
@@
$1/\sigma^{2}_{\bm{x(0)}}$ 2 1 1 0
DDPM $\zeta_{t}$ $\begin{cases}1\text{e-2}&\text{if }t>100\\ 1\text{e-3}&\text{otherwise }\end{cases}$ $\begin{cases}1\text{e-2}&\text{if }t>200\\ 1\text{e-3}&\text{otherwise }\end{cases}$ $\begin{cases}1\text{e-2}&\text{if }t>200\\ 1\text{e-3}&\text{otherwise }\end{cases}$ $\begin{cases}4\text{e-1}&\text{if }t>80\\ 4\text{e-4}&\text{otherwise }\end{cases}$
$1/\sigma_{\bm{x}(0)}^{2}$ 0.1 0.1 0.1 5e2
ImageNet DDIM $\zeta_{t}$ $\begin{cases}5\text{e-2}&\text{if }t>20\\ 1\text{e-2}&\text{otherwise }\end{cases}$ $\begin{cases}4\text{e-2}&\text{if }t>15\\ 8\text{e-3}&\text{otherwise }\end{cases}$ $\begin{cases}4\text{e-2}&\text{if }t>25\\ 1\text{e-2}&\text{otherwise }\end{cases}$ -
$1/\sigma_{\bm{x}(0)}^{2}$ 22 22 22 -
DDPM $\zeta_{t}$ $\begin{cases}2\text{e-2}&\text{if }t>200\\ 2\text{e-3}&\text{otherwise }\end{cases}$ $\begin{cases}1\text{e-2}&\text{if }t>200\\ 1\text{e-3}&\text{otherwise }\end{cases}$ $\begin{cases}5\text{e-3}&\text{if }t>100\\ 1\text{e-4}&\text{otherwise }\end{cases}$ $\begin{cases}2\text{e-2}&\text{if }t>80\\ 2\text{e-4}&\text{otherwise }\end{cases}$
$1/\sigma_{\bm{x}(0)}^{2}$ 1e21e-2 1e11e-1 11 $1\text{e1}$
$\times 4$ SR Gaussian Deblur Motion Deblur
$1/\sigma_{\bm{x}(0)}^{2}$ $\text{e-1} 11$ $1\text{e-1}$
Algorithm 2 CoDPS(DDPM)
0:NN, $\bm{y}$, $\zeta_{t}\}_{t=1}^{N},{\{\tilde{\sigma}_{t}\}_{t=1}^{N}}$, $\sigma_{x_{0}}^{2}$
@@
[…]
@@
DDS. We used the publicly available version of DDS [9]. We included our forward operator implementations, since the original code only supported medical imaging tasks. We tuned the regularizer parameter $\gamma$ between range 0.520.5-2 using grid search to obtain the best result for each problem.
@@
ReSample. We used the publicly available implementation of ReSample [16]. For each task, we used the corresponding configuration files. We modified the noise-level configuration to $\sigma_{y}=0.05$. For each problem, we tuned the scale parameters within the range 0.10.111.
@@ Appendix C Additional results and figures
Learning the data prior. In this experiment, we limit the number of mixtures to $k=3$. We build our dataset by generating samples from an $n$-dimensional GMM model, where $n$ is a randomly selected integer between $3$ and $15$. We randomly set the means of mixtures using random $n$ dimension vectors selected uniformly form the range [3,3][-3,3] and the covariance matrices using an $n\times n$ symmetric positive definite matrix whose entries are sampled uniformly from the range $[0,1]$. Now that we have a data at hand the first task it to learn its distribution using a diffusion model. We use an MLP-based model similar to the one used for toy problems in [2]. Our diffusion model consists of 4 MLP blocks, each with an input and output feature size of 64. Additionally, the model has input and output layers to map features between the data dimension $n$ and 64.
@@ C.2.1 Hyperparameter ablation study
To assess the sensitivity of our proposed methods to hyperparameters, we conducted an ablation study based on the reviewer’s suggestion. Specifically, we used the $\sigma_{0}^{2}$ and gradient scaling schedule $\zeta_{t}$} tuned for the super-resolution task (SR $\times 4$), and then applied the same hyperparameters directly to the Gaussian deblurring, motion deblurring, and inpainting tasks on the FFHQ dataset.
@@
Table 7: Ablation study using hyperparameters tuned for SR ($\times 4$) and applied directly to other tasks. Results show that CoDPS remains competitive without per-task tuning on two new tasks.
@@
SR ($\times 4$) Deblur (Gaussian) Deblur (motion)
@@
[…]
100 21.23 0.482 88.93 0.495 18.29 0.341 135.86 0.610 23.89 0.663 52.15 0.373
DPS [7]
1000 21.70 0.562 45.20 0.381 21.97 0.706 62.72 0.444 20.55 0.634 56.08 0.389
DiffPIR [6]
100 20.58 0.430 110.69 0.560 21.62 0.504 58.61 0.433 24.61 0.661 46.72 0.359
DDRM [5]
20 23.64 0.624 67.69 0.423 22.73 0.705 63.02 0.427 - - - -
MCG [20]
1000 13.39 0.227 144.5 0.637 16.32 0.441 95.04 0.550 5.89 0.037 186.9 0.758
PnP-ADMM [33]
- 20.43 0.538 159.58 0.592 20.77 0.545 105.96 0.502 21.02 0.563 108.53 0.500
ADMM-TV [34]
- 18.40 0.490 191.97 0.600 21.52 0.594 108.10 0.495 20.39 0.557 148.80 0.524
@@
Table 9: Performance metrics of inverse problems on ImageNet $\times256$-1k validation dataset.
@@
DPS [7]
25.02 0.750 32.21 0.269
DiffPIR [6]
DDRM [5]
22.06 0.593 84.44 0.437
MCG [20]
22.62 0.608 30.51 0.316
PnP-ADMM [33]
21.41 0.641 65.55 0.403
ADMM-TV
18.94 0.546 138.40 0.540